OpenStudy (anonymous):

Solve using L'Hôpital's rule: Lim x goes to 0 from the right , (cos x)^1/x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0+} (\cos x)^{1/x}\]

OpenStudy (math&ing001):

\[Ln ((\cos x)^{\frac{ 1 }{ x }})=\frac{ 1 }{ x }\ln(\cos x)\] Apply l'Hopital's rule on f(x)=ln(cos x) and g(x)=x to find the limit of the function above. Then use exponential to find your final answer.

OpenStudy (anonymous):

Thanks , so : if we tried the limit we'll get 0/0 and now we could use the rule , so: ln cosx/x = 1/cosx/x= 0 ?

OpenStudy (math&ing001):

No, you calculate \[\lim_{x \rightarrow 0^{+}}\frac{ f'(x) }{ g'(x) }\] L'Hôpital's rule stipulates that \[\lim_{x \rightarrow 0^{+}}\frac{ f'(x) }{ g'(x) }=\lim_{x \rightarrow 0^{+}}\frac{ f(x) }{ g(x) }=L\] When you find that limit, exp(L) will be your final answer.

OpenStudy (anonymous):

That's what I did , 1*ln(cosx)/x = ln (cosx)/x= the derivative of ( ln(cos x)) = 1/cox the derivative of x =1 so 1/cosx/1 = 1/1=1 ?

OpenStudy (math&ing001):

You have a problem on the derivative of ln(cos x)). \[(\ln (u))'=\frac{ u' }{ u }\]

OpenStudy (anonymous):

Thank you for being patient with me ^^ I used my notes and it says the derivative of ln x = 1/x anyways we'll : -sinx/cosx/1 so 0/1/1 = 0 ?

OpenStudy (math&ing001):

That's correct :) Now what would be the final answer?

OpenStudy (anonymous):

ln(0)=1

OpenStudy (math&ing001):

Yes, Gz!

OpenStudy (anonymous):

Thank you ^^

OpenStudy (math&ing001):

You're welcome ^^